Map deh needs to be given exactly 500 mL of a 15% (WY)intravenous glucose solton. The stock solution is55% (wV). How many mMters of the stock solution are needed to prepare the flu d forthe Ⅳbag? Which e1 how much water is needed to prepare the 15% (wh) glucose solution? O Volume given to patient (500 mL)/Volume of stock solution needed O Volume of stock solution needed /Volume given to patient (500 mL) O Volume of stock solution needed Volume given to patient (500 mL) O Volume given to patient (500 mL)- Volume of stock solution neededA patient needs to be given exactly 500 mL of a 15% (w/v) intravenous glucose solution. The stock solution is 55% (w/v). How many milliliters of the stock solution are needed to prepare the fluid for the IV bag?

A patient needs to be given exactly 500 mL of a 15% (w/v) intravenous glucose solution. The stock solution is 55% (w/v). How many milliliters of the stock solution are needed to prepare the fluid for the IV bag?
Answer #1

We can use the equation-

V1S1 = V2S2

V1 = ? The volume of stock solution,

S1 = 55% (w/v) is the strength.

V2 = 500 mL

S2 = 15% (w/v)

so, V1 = V2S2/S1 = 500*15/55 mL = 136.36 mL

136.36 mL of the stock solution are needed to prepare the fluid for the IV bag.